# Quasiparticle states of hexagonal BN: A van der Waals density functional study

### v1

@@ IV.4 Broader relevance of KI-CX usage
We consider KI-CX (which reduces to use of CX in KS-DFT when used for structure optimizations) as a faster alternative to GW and EOMCC studies when seeking to predict QPs and identify fingerprints of vdW interactions [RanPRB16]. The CX functional is, as a conserving vdW-DF [dion2004, berland2014, hybesc14, hyldgaard2020], designed to capture the vdW interactions directly within the MBPT framework, as far as possible. Of course, it may not capture all nonlocal-correlation effects and that is why continued testing (e.g., via KI-CX predictions) is important.
We repeat our observation that KI-CX(CX) stands out by proceeding with the atomic-structure and QP characterizations in concert. As stated in the introduction, the CX (in KS-DFT) and KI-CX (in KI-DFT) use exactly the same MBPT input to define the nonempirical XC functional, per se. We think that a possible success of the CX design as a dual-use (KS-DFT and KI-DFT) functional, is helped by its inherent emphasis on current conservation and thus, the implied reflection of the plasmon nature. On the one hand, the plasmon-energy shifts by electro-dynamical coupling [Berland_2015:van_waals, hyldgaard2020] is a natural framework to count the integrated effects of nonlocal-correlation effects [jerry65, hybesc14]. On the other hand, a consistent modeling of virtual currents (as in plasmons) are essential when moving electrons (as in the excitations that QPs characterize).
Our interest in KI-CX goes well beyond the present focus on testing. By use of KI-CX we may be able to provide fast predictions of atomic and electronic (QP) structure in concert, for example, tracking fingerprints of vdW and, in general, noncovalent interactions in DNA systems while comparing with, e.g., optimally-tuned RSH vdW-DFs [JPCM2025].
@@ Acknowledgements.
$c$ [Å] $N^{\|}_{k}$ $E_{\rm cut}^{\rm wf}$ [Ry] $\text{G}_{v}\rightarrow\text{G}_{c}$ $\text{M}_{v}\rightarrow\text{M}_{c}$ $\text{K}_{v}\rightarrow\text{G}_{c}$ $\text{K}_{v}\rightarrow\text{K}_{c}$ $\text{M}_{v}\rightarrow\text{G}_{c}$ $\text{K}_{v}\rightarrow\text{M}_{c}$
20 7 100 7.62 8.98 5.90 7.94 6.79 8.08
20 7 200 8.14 9.90 6.42 8.61 7.31 9.00
20 7 300 8.18 9.86 6.46 8.62 7.36 8.96
20 7 400 8.21 9.92 6.48 8.69 7.38 9.02
30 7 100 8.36 10.01 6.63 8.81 7.53 9.11
30 7 200 8.28 9.95 6.55 8.65 7.45 9.05
30 7 300 8.28 10.17 6.56 8.89 7.45 9.27
30 7 400 8.20 9.99 6.47 8.72 7.37 9.09
20 7 200 8.14 9.90 6.42 8.61 7.31 9.00
20 9 200 7.70 8.90 5.94 8.01 6.85 7.99
20 11 200 8.08 9.59 6.32 8.40 7.23 8.67
20 7 400 8.21 9.92 6.48 8.69 7.38 9.02
20 9 400 8.34 9.58 6.58 8.75 7.49 8.66
20 11 400 7.69 9.15 5.93 7.93 6.84 8.24
Table 4: Search for precision (numerical stability with respect to choice of modeling parameters) in KI-CX predictions of BN1 QP-gap energies (in eV) as defined by transitions between indicated valence- and conduction-band symmetry points. Italicized table entries identify the set of model parameters that we use, in the main text, to report and discuss the QP behavior for the strictly two-dimensional BN1 system.
$E_{\rm cut}^{\rm wf}$ [Ry] $\text{G}_{v}\rightarrow\text{G}_{c}$ $\text{M}_{v}\rightarrow\text{M}_{c}$ $\text{K}_{v}\rightarrow\text{G}_{c}$ $\text{K}_{v}\rightarrow\text{K}_{c}$ $\text{M}_{v}\rightarrow\text{G}_{c}$ $\text{K}_{v}\rightarrow\text{M}_{c}$ Fundamental Gap
150a 9.14 7.06 7.50 7.46 7.85 6.71 6.58b
200a 9.14 7.06 7.50 7.46 7.85 6.71 6.58b
250a 9.14 7.06 7.50 7.46 7.85 6.71 6.58b
300a 9.14 7.06 7.50 7.46 7.85 6.71 6.59b
350a 8.84 7.41 7.20 7.46 7.55 7.07 6.93b
$GW^{c}$ 8.40 6.50 6.90 6.90 7.30 - 5.95b
Exper.d - - - - - - 6.08
a KI-CX with $7\times 7\times 3$ $k$-point sampling.
b Corresponds to $\text{T}_{v}\rightarrow\text{M}_{c}$.
c Ref. hunt2020.
d Ref. arnaud2006.
Table 5: FOR here but not in archive and submitted versions unless surprise with double checks. QP gap predictions (in eV) by KI-CX for h-BN bulk structure $A$ as a function of the choice of the choice of the wavefunction-energy cut off $E_{\rm cut}^{\rm wf}$. The highlighted parameter choice identify the pick we made for discussing the nature and and energy variation of QPs in bulk-h-BN characterization in the main text.
@@ Appendix A Linear response for nonlocal correlations
As mentioned in Section II.B, the DFPT response of the non-local potential $v^{\text{nl}}(\mathbf{r})$ is implemented in the kcw module of QE. The necessary linear-response steps are derived in Ref. sabatini2016 with respect to a general perturbation $\Delta^{\lambda}v^{\text{nl}}(\mathbf{r})$ and with the purpose (there) of computing vdW signatures on phonons and their interactions. Here, we summarize how this DFPT formulation for $\Delta^{\lambda}v^{\text{nl}}(\mathbf{r})$ appears when considered with respect to the removal or addition of a fractional electron, i.e., for KI-DFT use.
As a starting point, consider the definition of the non-local energy in Eq. (6), approximating the spatial integrals by sums over a uniform grid of points (at fixed positions $\rho_{i}$ and $\mathbf{r}_{i}$) and the gradients by finite differences on the same grid. This allows the use of partial derivatives of the density variation on the grid to be used in place of formal functional derivatives [roso09]. Applying these finite differences and simplifying, we obtain:
@@
where $b_{i}$ corresponds a term defined exclusively by differentiating direct by density (at the specific grid site $i$). The second term, defined by a set of $h_{j}$ contributions, reflects instead derivations with respect to the density gradients. The evaluation of a gradient in QE makes use of specific stencil so that the gradient at site $i$ can be expressed in terms of the density values at neighboring grid sites $j$ [roso09].
@@
Expressing also the $v^{\text{nl}}_{i}$ for in Ref. (16) reveals the second derivatives present in Eq. (10). To get a numerical evaluation, we adapt the grid-specific implementation [roso09, sabatini2016, giannozzi2017] of the formal determination
@@
The details of these two components are explicitly derived, for the stencil used in QE, in Ref. sabatini2016.
@@ Appendix B Precision in KI-CX
For KS-DFT studies using norm-conserving pseudopotentials, past experience suggest that a choice $E_{\rm cut}^{\rm wf}=200$ Ry for wavefunction kinetic-energy cut off is sufficient to converge predictions of total-energy differences and structure [shukla2022, JPCM2025].
@@
For descriptions of conduction-band QP states we see a need to secure numerical stability with regards to changes in the choice of implicit modeling parameters; This observation has practical implications for our KI-CX study of the single-sheet BN1 system. The description of unoccupied QPs is generally a challenge for KS-DFT, for example, documented for electron attachment in nitrogen-base systems [nguyen2016, JPCM2025]. This problem emerges, in part, because the spatial variation of unoccupied KS-orbitals does not naturally enter in the iterative optimization of the actual ground-state electron density $\rho(\mathbf{r})$. In contrast, with a hybrid like AHBR-mRSH and with the KI-DFT framework, we can overcome also this KS-DFT problem to accurately describe the empty QPs [nguyen2016, JPCM2025] because we make the iterative computational scheme dependent on the nature of these empty orbitals. However, the requirements on the $E_{\rm cut}^{\rm wf}$ choice for stability in the description of conduction-band states might well be higher. Also, given our screening handling [gygi1986], we must validate that the chosen height $c$ of the modeling cell for the strictly two-dimensional BN1 system is sufficient to electrically decouple the sheets.
Table 4 presents an overview of the numerical work we completed to assert the level of precision we presently have for a KI-DFT description of (empty-state) QP energies in BN1. We see that the step of crafting of the MLWF representation for KI-DFT use can sometimes cause abrupt changes in predictions of conduction-band QP energies. A complication is that the $E_{\rm cut}^{\rm wf}$ variation also depends on the choice of $c$. Our criteria for reporting BN1 QP predictions is that we reach a fair numerical stability with regards to changes in both of these modeling parameters. Tables 2 and 4 report our finding that we have a fair level of cross-checked stability for the conduction-band QPs of BN1 when using $c=20$ Å and $E_{\rm cut}^{\rm wf}=400$ Ry, as also stated in the main text.
NOTE: The choice of $N_{\|}$ in a set of possible $N_{k}^{\|}\times N_{k}^{\|}\times 1$ samplings of $k$ points for in BN1 predictions has also been tested. If the choice of $N_{k}^{\|}=7$ is sufficient to reflect a plateau of BN1 predictions also in the modeling aspect, as is suggested in Table 4, then we shall merely state that as a comment here in the appendix text.
NOTE: Table 5 summarizes a study of numerical stability for the $A$ bulk system. This table will not be in arXiv or submitted version BUT it is convenient to have here for overview in house.
We also performed a study of numerical stability for the bulk $A$ system in KI-CX. We found no variation in predictions for QP gap results as we increased $E_{\rm cut}^{\rm wf}$ from 150 Ry to 300 Ry. Accordingly we chose to complete all KI-DFT characterizations for bulk h-BN at the value $E_{\rm cut}^{\rm wf}=200$ Ry that is adequate for total-energy and structure characterizations in KS-DFT.

### v2

@@ IV.4 Broader relevance of KI-CX usage
We consider KI-CX (which reduces to use of CX in KS-DFT when used for structure optimizations) as a faster alternative to GW and EOMCC studies when seeking to predict QPs and identify fingerprints of vdW interactions [3]. The CX functional is, as a conserving vdW-DF [70, 78, 71, 35], designed to capture the vdW interactions directly within the MBPT framework, as far as possible. Of course, it may not capture all nonlocal-correlation effects and that is why continued testing (e.g., via KI-CX predictions) is important.
We repeat our observation that KI-CX(CX) stands out by proceeding with the atomic-structure and QP characterizations in concert. As stated in the introduction, the CX (in KS-DFT) and KI-CX (in KI-DFT) use exactly the same MBPT input to define the nonempirical XC functional, per se. We think that a possible success of the CX design as a dual-use (KS-DFT and KI-DFT) functional, is helped by its inherent emphasis on current conservation and thus, the implied reflection of the plasmon nature. On the one hand, the plasmon-energy shifts by electro-dynamical coupling [72, 35] is a natural framework to count the integrated effects of nonlocal-correlation effects [95, 71]. On the other hand, a consistent modeling of virtual currents (as in plasmons) are essential when moving electrons (as in the excitations that QPs characterize).
Our interest in KI-CX goes well beyond the present focus on testing. By use of KI-CX we may be able to provide fast predictions of atomic and electronic (QP) structure in concert, for example, tracking fingerprints of vdW and, in general, noncovalent interactions in DNA systems while comparing with, e.g., optimally-tuned RSH vdW-DFs [23].
@@ Acknowledgements.
$c$ [Å] $E_{\rm cut}^{\rm wf}$ [Ry] $\text{G}_{v}\rightarrow\text{G}_{c}$ $\text{M}_{v}\rightarrow\text{M}_{c}$ $\text{K}_{v}\rightarrow\text{G}_{c}$ $\text{K}_{v}\rightarrow\text{K}_{c}$ $\text{M}_{v}\rightarrow\text{G}_{c}$ $\text{K}_{v}\rightarrow\text{M}_{c}$
20 100 7.62 8.98 5.90 7.94 6.79 8.08
20 200 8.14 9.90 6.42 8.61 7.31 9.00
20 300 8.18 9.86 6.46 8.62 7.36 8.96
20 400 8.21 9.92 6.48 8.69 7.38 9.02
30 100 8.36 10.01 6.63 8.81 7.53 9.11
30 200 8.28 9.95 6.55 8.65 7.45 9.05
30 300 8.28 10.17 6.56 8.89 7.45 9.27
30 400 8.20 9.99 6.47 8.72 7.37 9.09
Table 4: Search for precision (numerical stability with respect to choice of modeling parameters) in KI-CX predictions of BN1 QP-gap energies (in eV) as defined by transitions between indicated valence- and conduction-band symmetry points. Italicized table entries identify the set of model parameters (along with corresponding gap results) that we use, in the main text, to report and discuss the QP behavior for the strictly two-dimensional BN1 system.
@@ Appendix A Linear response for nonlocal correlations
As mentioned in Section II.B, the DFPT response of the non-local potential $v^{\text{nl}}(\mathbf{r})$ is implemented in the kcw module of QE. The necessary linear-response steps are derived in Ref. 117 with respect to a general perturbation $\Delta^{\lambda}v^{\text{nl}}(\mathbf{r})$ and with the purpose (there) of computing vdW signatures on phonons and their interactions. Here, we summarize how this DFPT formulation for $\Delta^{\lambda}v^{\text{nl}}(\mathbf{r})$ appears when considered with respect to the removal or addition of a fractional electron, i.e., for KI-DFT use.
As a starting point, consider the definition of the non-local energy in Eq. (6), approximating the spatial integrals by sums over a uniform grid of points (at fixed positions $\rho_{i}$ and $\mathbf{r}_{i}$) and the gradients by finite differences on the same grid. This allows the use of partial derivatives of the density variation on the grid to be used in place of formal functional derivatives [96]. Applying these finite differences and simplifying, we obtain:
@@
where $b_{i}$ corresponds a term defined exclusively by differentiating direct by density (at the specific grid site $i$). The second term, defined by a set of $h_{j}$ contributions, reflects instead derivations with respect to the density gradients. The evaluation of a gradient in QE makes use of specific stencil so that the gradient at site $i$ can be expressed in terms of the density values at neighboring grid sites $j$ [96].
@@
Expressing also the $v^{\text{nl}}_{i}$ for in Ref. (16) reveals the second derivatives present in Eq. (10). To get a numerical evaluation, we adapt the grid-specific implementation [96, 117, 81] of the formal determination
@@
The details of these two components are explicitly derived, for the stencil used in QE, in Ref. 117.
@@ Appendix B Precision in KI-CX
For KS-DFT studies using norm-conserving pseudopotentials, past experience suggest that a choice $E_{\rm cut}^{\rm wf}=200$ Ry for wavefunction kinetic-energy cut off is sufficient to converge predictions of total-energy differences and structure [74, 23].
@@
For descriptions of conduction-band QP states we see a need to seek numerical stability with regards to changes in choices for implicit modeling parameters; This observation has practical implications for our KI-CX study of the single-sheet BN1 system. The description of unoccupied QPs is generally a challenge for KS-DFT, for example, documented for electron attachment in nitrogen-base systems [60, 23]. This problem emerges, in part, because the spatial variation of unoccupied KS-orbitals does not naturally enter in the iterative optimization of the actual ground-state electron density $\rho(\mathbf{r})$. In contrast, with a hybrid like AHBR-mRSH and with the KI-DFT framework, we can overcome also this KS-DFT problem to accurately describe the empty QPs [60, 23] because we make the iterative computational scheme dependent on the nature of these empty orbitals. However, the requirements on the $E_{\rm cut}^{\rm wf}$ choice for stability in the description of conduction-band states might well be higher. Also, given our screening handling [125], we must validate that the chosen height $c$ of the modeling cell for the strictly two-dimensional BN1 system is sufficient to electrically decouple the sheets.
Table 4 presents an overview of the numerical work we completed to assert the level of precision we presentlty have for a KI-DFT description of (empty-state) QP energies in BN1. We see that the step of crafting the MLWF representation can sometimes cause abrupt changes in the resulting predictions of conduction-band QP energies. A complication is that the $E_{\rm cut}^{\rm wf}$ variation also depends on the choice of $c$. Our criterion for using these BN1 QP predictions is that we reach a fair numerical stability with regards to changes in both these modeling parameters. Tables 2 and 4 report our finding that we have a fair cross-checked stability for the conduction-band QPs of BN1 when using $c=20$ Å and $E_{\rm cut}^{\rm wf}=400$ Ry, as also stated in the main text.
We also performed a study of numerical stability for the bulk $A$ system in KI-CX but found no variation in predictions for QP gap results as we increased $E_{\rm cut}^{\rm wf}$ from 150 Ry to 300 Ry. Accordingly we choose to complete all KI-DFT characterizations for bulk h-BN at the value $E_{\rm cut}^{\rm wf}=200$ Ry that is adequate for total-energy and structure characterizations in KS-DFT.